What is the budget for a processing unit that is needed by the city perpetually if it needs an initial investment of $5 million and an investment of $5 million every 20 years? The rate the city uses is 3% p y c y.
Answer #1

Capitalized Investment=5+5/1.03^20*1/(1-1/1.03^20)=11.20261793 million

Annual Investment=(5+5/1.03^20*1/(1-1/1.03^20))*3%=0.33607854 million